Dan is pedaling at a resistance (R) of 1.0kg on a Monarkarm ergometer. He is pedaling at 50 rev/min and weight 90kg. Determine their Caloric Cost (kcal/min).
Answer #1

Distance , D = Cadence(rev/min)*constant(6m)

D = 50rev/min*6m = 300m/min

Pedalling resistance, R = 1.0 kg

Work rate, W = R*D = 1.0*300 = 300 kgm/min

Arm VO2 = 3(work rate)/body mass+3.5 = 3*300/90 +3.5 = 13.5 ml*kg-1*min-1

Absolute VO2 = 13.5*90 ml*min-1 = 1215 ml*min-1=1215/1000 L*min-1 = 1.215 L*min-1

Caloric cost = 1.215*5 Kcal/min = 6.075 Kcal/min